OpenStudy (anonymous):

Juliet has attempted 213 problems on Brilliant and solved 210 of them correctly. Her friend Romeo has just joined Brilliant, and has attempted 4 problems and solved 2 correctly. From now on, Juliet and Romeo will attempt all the same new problems. Find the minimum number of problems they must attempt such that it is possible that Romeo's ratio of correct solutions to attempted problems will be strictly greater than Juliet's.

hero (hero):

I remember this question. Let's see if I can remember how to set it up.

hero (hero):

Hang on...let me check something. BRB

hero (hero):

Okay, I have it.

hero (hero):

Juliet's ratio: \[\frac{210}{213}\] Romeo's ratio: \[\frac{2}{4}\] Basically, the only way Romeo's ratio beats Juliet is if Juliet stops getting answers right In other words: \[\frac{2 + x}{4 + x} > \frac{210}{213 + x}\]

hero (hero):

Solve that for x

hero (hero):

Oh and \(x\) equals the minimum number of problems they will attempt

hero (hero):

Probably should be \(\ge\) sign

hero (hero):

This is the correct setup. All you need to do is solve for \(x\). How is that going so far @marb0i

hero (hero):

Your solution will obviously involve a positive number. Discard any negative results.

OpenStudy (anonymous):

Is the answer a whole number? Beacause mine contains a decimal.

hero (hero):

Yes, the correct answer is a whole number. Mind sharing your steps? I can point you in the right direction.

hero (hero):

And then, what happened after that? You should post all your steps.

OpenStudy (anonymous):

\[\frac{ 2+x }{ 4+x } \ge \frac{ 210 }{ 213 + x }\] \[(2 + x)(213 + x) = (4+x)(210)\] \[426 + 2x + 213x +2x = 840 + 210x\] \[426 - 840 = 210x - 217x\] x =138 Am i doing this right? sorry my net is so crappy

hero (hero):

No that is not correct. You have multiplied wrong

hero (hero):

There should be an \(x^2\) in there somewhere.

hero (hero):

\(426 + 2x + 213x +x^2 \ge 840 + 210x\)

hero (hero):

That should be the correct inequality after cross multiplication